PROBLEM 4.2 (pg 63, #17) Let A denote the event that the next request for assistance from a statistical software consultant relates to the SPSS package, and let B denote the event that the next request is for help with SAS. Suppose that P(A) -0.25 and P(B)-0.65. Assume that any request for help never relates to two or more packages simultaneously. Note: since P -0.9, which is less than 1, there are clearly other packages for which help is requested. (A)+P(B) P(A P(AUB) a. Calculate b. Calculate P((AUB) cCalct) which equis (4UB)) which equals
Answer #1

SolutionA:

P(Ac)=1-P(A)

=1-0.25

=0.75

ANSWER:0.75

Solutionb:

P(AUB)=P(A)+P(B)-P(A and B)

P( A and B)=P(A)*P(B) since both are independent events since outcome of one event does not effect the outcome of other

therefore

P(AUB)=0.25+0.65-0.25*0.65

P(AUB)=0.7375

ANSWER:0.7375

Solutionc:

P(AUB)c= 1-P(AUB)

=1-0.7375

=0.2625

ANSWER:

0.2625